OpenStudy (anonymous):

solve for n: 3n+2/7 = 1- 2n-3/4 -1/26 41/26 10/13 7/13

OpenStudy (anonymous):

Please put the required parts in brackets eg; 3n + (2/7) is different from 3n+2/7. Please write the qn that way then i can do it.

OpenStudy (anonymous):

3n+2 2n-3 ----- =1 ------- 7 4

OpenStudy (anonymous):

ok. We have - (3n+2)/7 = [4-(2n-3)] / 4 => 12n+8 = 7(4+3-2n) 12n+8 = 7(7-2n) 12n+8 = 49-14n 26n = 41 N = 41/26